Which of the following bases can deprotonate a terminal alkyne (check all that apply)? CH3CH2Li CH3CH2ON NaH NaNH2 t-BuOK
Answer #1

phoratue of BuOH is 16.54 at 25°c HC CH₂OH is = 15.9 at 25°c HC CH₂ is = 50 at 25c. NH₂ is = 38 at 25€ R-C=C-H is = 26 at 25c